Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $110 per unit.
Answer #1
Computation of Contribution margin per unit:-
Sales price per unit $110
Less: Variable cost per unit $77
Contribution margin per unit $33
Fixed Costs per month $161,700
1 Computation of Company's Break-even point in unit sales:-
Break even point in units = Total Fixed Costs
Contribution per unit
= $161,700
$33
= 4900
Computation of Company's Break-even point in dollar sales :-
a Break even point in units 4900
b Selling Price Per unit $                                110
c Break-even point in dollar sales (a*b) $                        539,000
2 If variable expenses per stove increase as a % of the selling price, will it result in higher or lower break even point?
It will result in higher break even point because due to increase in variable expenses, contribution margin ratio get reduced and the it results into higher break even point.
3 Contribution Format Income Statements:-
Proposed Production and sales = 19000 + 19000*25% = 23750 stoves
Reduced Selling Price = 110 - 10% of 110 = $99 per stove
Outback Outfitters
Contribution Income Statement
Present Proposed
Production Capacity (stoves) 19000 stoves 23750 stoves
Per Unit Total Per Unit Total
a Sales $                                110 $                  2,090,000 $                               99 $   2,351,250
b Variable cost $                                  77 $                  1,463,000 $                               77 $   1,828,750
c Contribution margin (a-b) $                                  33 $                     627,000 $                               22 $      522,500
d Fixed costs $                     161,700 $       161,700
e Net Operating Income (c-d) $                     465,300 $      360,800
4. Computation of No. of stoves to be sold at new selling price to attain a target profit of $72000 per month:-
Computation of Revised contribution margin :-
Sales price per unit $                               99
Less: Variable cost per unit $                               77
Contribution margin per unit $                               22
a Fixed Expenses $                        161,700
b Target Profit $                           72,000
c Desired Contribution (a+b) $                        233,700
d Contribution margin per unit $                                  22
Unit sales needed to attain target profit (Sales Volume) = Desired Contribution
Contribution per unit
= $233,700
$22
= 10623 stoves

Feel free to ask any clarification, if required. Kindly provide feedback by thumbs up, if satisfied. It will be highly appreciated.
Thank you.